Circular No. 7046

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998eo, 1998ep, 1998eq G. Aldering, Lawrence Berkeley National Laboratory, on behalf of the

Supernova Cosmology Project (cf. IAUC 6804), reports the discovery of three type-Ia supernovae, all found as part of their high-redshift supernova search on the Keck II telescope:

SN  1998 UT  α_{2000}  δ_{2000}  I z  Offset
1998eo Oct. 15 4^{h}56^{m}15.^{s}47 − 3^{o}46′38′′.7 23.6 0.84 34′′ W, 24′′ S
1998ep Oct. 15 23 19 57.67 +16 03 19.5 20.1 0.11 near center
1998eq Oct. 15 23 20 27.47 +15 55 43.7 24.8 1.20 near center

The supernova spectra were obtained at the Keck II telescope on Oct. 27. The redshifts were determined from host-galaxy emission and absorption lines. Both SN 1998eo and SN 1998eq are within 1′′ of the host-galaxy center. I-band photometry of the lightcurves is being obtained from the ground for all supernovae with determined redshifts, and SN 1998eq and SN 1998eo have been observed in the I band with the Hubble Space Telescope; J-band photometry for SN 1998eo has been obtained with NICMOS on the HST. The spectrum of SN 1998ep indicates that it was roughly two weeks after maximum on Oct. 27, while the spectra of SN 1998eo and SN 1998eq indicate they were near maximum.

SUPERNOVA 1998er IN GH 9-2 A. Becker, University of Washington, on behalf of the MACHO Col-

laboration, reports the discovery of an apparent supernova, possibly before maximum brightness, in the LMC background galaxy GH 9-2: “The object is question was initally announced as an on-going microlensing event (MACHO Alert 98-LMC-4) on an object located at α = 5^{h}50^{m}16^{s}.87, δ = −71^{o}14′13′′.5 (equinox 2000.0). Inspection of the images shows that this object is superimposed on the outer visible edge of the background galaxy GH 9-2, identified in Gurwell and Hodge (1990, PASP 102, 849), and is therefore likely a supernova. The morphology of GH 9-2 suggests that this is a late-type SBb galaxy, and SN 1998er is located 10′′.5 west and 4′′.8 north of the galaxy nucleus. Images confirming this object’s brightening were taken with the Cerro Tololo 0.9-m telescope on Nov. 3.346 and 4.342 UT. As of Nov. 4.491, the object was at magnitudes R = 17.4 and V = 18.6. A finding chart and additional information are available by anonymous ftp

at darkstar.astro.washington.edu:/macho/Alerts/98-LMC-4 or at the
WWW site http://darkstar.astro.washington.edu.”

1998 November 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green